5. Dr. Wong orders 20 grams of a 52% solution of a certain medicine. However, the
pharmacist only has bottles of 40% and 70% solution. How much of each should the
pharmacist mix to Mr. Wong's satisfaction?

Answer by venugopalramana(3286)

5. Dr. Wong orders 20 grams of a 52% solution of a certain medicine. However, the
pharmacist only has bottles of 40% and 70% solution. How much of each should the
pharmacist mix to Mr. Wong's satisfaction?
LET X GMS OF 40% SOLUTION BE MIXED WITH 20-X GMS OF 70% SOLUTION.
MEDICINE CONTENT IN MIX = X*40/100 + (20-X)*70/100=20*52/100
0.4X+14-0.7X=10.4
0.7X-0.4X=14-10.4=3.6
0.3X=3.6
X=3.6/0.3=12 GMS OF 40% SOLUTION AND 20-12=8 GMS OF 70% SOLUTION ARE TO BE MIXED TO GET 20 GMS. OF 52% SOLUTION.